Given 𝑑𝑦





𝑑𝑥





= 𝑥 𝑥





2 + 𝑦





2





𝑒





−𝑥





and 𝑦(0) = 1, 𝑦(0.1) = 1.005, find





i) 𝑦 0.2 by Euler’s method





ii) 𝑦 0.3 by Modified Euler’s method





iii) 𝑦(0.4) by Milne’s predictor - corrector method.

1
Expert's answer
2022-02-22T15:23:16-0500

"\frac{dy}{dx}=x^2+y^2e^{-x}=f(x,y)\\"

i)

By Euler's method,

"y_{n+1}=y_n+hf(x_n,y_n)\\
y_2=y_1+0.1×[x_1^2+y_1^2e^{-x_1}]\\
=1.005+0.1×[0.1^2+1.005^2e^{-0.1}]\\
=1.005+0.1×0.9239
=1.0973"


ii)

By Euler's method,

"y_{n+1}=y_n+hf(x_n,y_n)\\
y_3=y_2+0.1×[x_2^2+y_2^2e^{-x_2}]\\
=1.0973+0.1×[0.2^2+1.0973^2e^{-0.2}]\\
=1.0973+0.1×0.9858
=1.1958"


iii)

By Milne’s predictor method,

"y_{n+1,p}=y_{n-3}+\frac{4h}{3}(2f_{n-2}-f_{n-1}+2f_n)\\
y_{4,p}=y_{0}+\frac{4×0.1}{3}(2f_{1}-f_{2}+2f_3)\\
=1+\frac{0.4}{3}(2×1.005-1.0973+2×1.1958)\\
=1+0.4405\\
=1.4405"


By Milne’s corrector method,

"y_{n+1,c}=y_{n-1}+\frac{h}{3}(f_{n+1}+f_{n-1}+4f_n)\\
y_{4,p}=y_{2}+\frac{0.1}{3}(f_{4}+f_{2}+4f_3)\\
=1.0973+\frac{0.1}{3}(1.4405+1.0973+4×1.1958)\\
=1.0973+2.1963\\
=3.2936"